The inventory records of Martin Corporation reflected the following information for the month of August:

DateTransactionNumber of UnitsUnit Cost

8/1Beginning inventory400$5

8/3Purchase No. 1400$5

8/5Sale No. 1600

8/7Sale No. 2100

8/11Purchase No. 21,000$7

8/17Sale No. 3700

8/19Purchase No. 31,000$7

8/21Sale No. 4600

8/28Sale No. 5600

8/29Purchase No. 41,200$9

8/30Ending inventory


Required:

A. Determine the amount of the ending inventory and cost of goods sold under each of the following methods assuming the periodic inventory system.


MethodEnding
Inventory
Cost of Goods Sold

a.Average cost$$

b.FIFO$$

c.LIFO$$


B. Why would cash flow considerations relate to the choice of an inventory method?
Answer #1

Part 1

Method

Ending inventory

Cost of goods sold

a. Average cost

$10080

$18720

b. FIFO

$12200

$16600

c. LIFO

$8200

$20600

In FIFO Method, inventory is issued first from first purchase, then second purchase and so on

In LIFO method, inventory is issued first from last purchase, then second last purchased and so on

In average cost method, inventory is issued at average cost

FIFO

Number of units available for sale = 400+400+1000+1000+1200 = 4000

Units sold = 600+100+700+600+600 = 2600

Units in ending inventory = 4000-2600 = 1400

Cost of goods sold = (400*5)+(400*5)+(1000*7)+(800*7) = $16600

Ending inventory = (200*7)+(1200*9) = $12200

LIFO

Cost of goods sold = (1200*9)+(1000*7)+(400*7)= $20600

Ending inventory = (400*5)+(400*5)+(600*7) = $8200

Average method

Average cost = total cost / total units = ((400*5)+(400*5)+(1000*7)+(1000*7)+(1200*9))/(400+400+1000+1000+1200) = 7.20

Cost of goods sold = 2600*7.20 = $18720

Ending inventory = 1400*7.20 = $10080

Part 2

The cost of inventory method used by the company has significant impact on the cash flows. During the times of inflation or rising prices, the cost of goods sold is lower in case of FIFO while higher in case of LIFO method. This have impact on income and finally to tax liability. The tax payment can be reduced during the inflation by making use of LIFO method.